What is the square root of 5/8? (and how do you get the answer, please)
Answer by edjones(7569)

sqrt(5)/sqrt(8)
=sqrt(5)/sqrt(8) * sqrt(8)/sqrt(8) to rationalize the denominator
=sqrt(40)/sqrt(64)
=sqrt(10*4)/8
=2sqrt(10)/8
=sqrt(10)/4
.
Ed